OpenStudy (anonymous):

Solve the system of equations.3x - 4y = -10 and 5x + 8y = -2

OpenStudy (anonymous):

My first answer was (-2,4) it is really (-2,1)

OpenStudy (anonymous):

Answer: (-2,1) 3x-4y=-10 5x+8y=-2 ----------- 6x-8y=-20 multiply each number by 2 in order to cancel out 5x+8y=-2 ----------- 11x+0=-22 x=-2 plug in for y.